20 years hence, a man will be 3 times as old as he was 20 years ago. His present age is?
Answer by TimothyLamb(4379)

x + 20 = 3(x - 20)
x + 20 = 3x - 60
2x = 80
x = 40
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php